OpenStudy (anonymous):

solve for x & simplify as much as possible 6x-34 = -2 (x-3)

OpenStudy (anonymous):

6x-34 = -2(x-3) Distribute 6x-34=-2x+6 Then combine like terms, add 2x to both sides 2x 2x 8x-34=6 Add 34 to both sides 8x=40 Divide both sides by 8 x=5

OpenStudy (anonymous):

@bumblebees01234 do u understand ?

OpenStudy (anonymous):

yes. thanks. could you do a couple more?? please

OpenStudy (anonymous):

Fire-away

OpenStudy (anonymous):

-5 (y+9) = 3y-21

OpenStudy (anonymous):

-5y-45 =3y-21 Distribute -45=8y-21 Add 5y to both sides -24 = 8y Add 21 to both sides y=-3 Divide both sides by 8

OpenStudy (anonymous):

5y + 20 = -7(y-8) last one

OpenStudy (anonymous):

5y+ 20=-7y +56 Add -7 to both sides 12y+20= 56 Subtract 20 12y=36 y=3

OpenStudy (anonymous):

thank you so much! can you calssify parallelograms??

OpenStudy (anonymous):

close this q and open a new one :)